Your mother is using an electric flat iron that draws a current of 15 A when connected to a 220 V line.

(a) What is the resistance of the heating element of the flat iron assuming that it obeys Ohm’s law? What is the resistivity of the heating element if it is a wire that is 2.7 cm long with a cross-sectional area of 2.1 x 10-4 m2?



1
Expert's answer
2022-02-28T10:41:21-0500

(a)

"R=\frac{V}{I}=\frac{220}{15}=14.7\:\Omega"

(b)

"\rho=\frac{RA}{l}=\frac{14.7*2.1*10^{-4}}{0.027}=0.11\:\rm \Omega\cdot m"